Operations Report – February 9th

Crew 188 Operational Report 09/02/2018

Sol 12

Zac Trolley

Non-Nominal Systems: electrical system, electrical cords, propane, EVA Suits

Notes on non-nominal systems: See individual systems at the end of the report.

Generator (hours run): 14.5h

Solar: SOC 80%

Diesel: 72%

Propane: 55

Ethanol Free Gasoline (5 Gallon containers for ATV): 3 Gallons

Water (trailer): 0 Gallons

Water (static): 65 Gallons

Trailer to Static Pump used: No

Water (loft) – Static to Loft Pump used: Yes

Meter: 130949.8

Toilet tank emptied: No

ATVs Used: Yes

Oil Added: No

ATV Fuel Used: 0 Gallons

# Hours the ATVs were Used today: 0.5

Notes on ATVs: The Gasoline storage area needs to be moved. The sun is creating a large pressure differential on the jerry cans, making that area an explosion risk.

Deimos rover used: No

Hours: Hours: 108.1

Beginning charge:

Ending charge:

Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: No

Hours: 19.5h

Beginning charge:

Ending charge:

Currently charging: No

Opportunity rover used: No

Hours: 19.1h

Beginning charge:

Ending charge:

Currently charging: No

Curiosity rover used: No

Hours: 12.0h

Beginning charge:

Ending charge:

Currently charging: No

HabCar: No

General notes and comments:

During an EVA, a plastic handle from the passenger side roll cage on Demos came off. We have the handle, but no hardware to re-attach it.

Generator was turned off at 9:25

Generator was turned on at 18:24

I’ve found two extension cords that have been damaged and need to be replaced. If there are 3 prong plugs available, I can repair the cords. Please advise.

Summary of internet: Tests were run on the internet connectivity and passed on to Bernard.

Summary of suits and radios: I’ve passed on a report about the EVA suits

Summary of Hab operations: Propane was smelt again this afternoon around 4pm. The heater had been turned off for 24 hours, and nothing else was on. I strongly suggest a licensed contractor investigate the propane leak before another team moves into Hab. There is a significant safety risk as we have not been able to pinpoint the source of the leak.

Summary of EVA operations: Nominal

Summary of GreenHab operations: The temperature gauge display has some non-functional LCD areas, a new one has been ordered.

Summary of Science Dome operations: The science dome is awaiting a portable heater as per the MDRS Director.

Summary of RAM operations: Not Operational.

Summary of health and safety issues: Nominal

Questions, concerns and requests to Mission Support: We could use another bag of toilet deodorizer, as we only have one puck left.

Copyright © The Mars Society. All rights reserved. | Main Site